# Jets are the most robust observable ingredient of common envelope evolution

### v1

@@ 1 Introduction
There are more than a hundred bipolar and elliptical PNe, including some post-asymptotic giant branch nebulae and pre-PNe, having a central binary system (e.g., Miszalski et al. 2019c; Orosz et al. 2017; Jones 2020, 2024). The short orbital periods of these binary systems show that they have experienced a common envelope evolution (CEE) phase. In post-AGB systems and PNe with long orbital periods, months to years, the systems might have experienced the grazing envelope evolution (GEE). The morphologies of many of these suggest that jets shaped these PNe, pointing at the direct link between CEE and jet-launching. In this study, I quantify this occurrence rate and compare it to the occurrence rate of equatorial outflows (ring/torus/disk).
@@ 2 The sample of post-CEE PNe with jets
I examined the catalog of PNe with a binary central star built and maintained by David Jones [^1] https://www.drdjones.net/bCSPN/ (Jones & Boffin, 2017; Boffin & Jones, 2019; Jones, 2024). I do not consider the uncertain cases in that catalog (marked by * or ** there). This leaves 108 PNe. I could find high-quality images to decide whether jets exist in 50 PNe. Out of these 50 PNe, I saw clear indications for jets or morphologies robustly suggesting jets in 28 PNe, marked by the number ‘1’ in the third column of Table 1. In 8 PNe, I found morphologies that likely suggest jets, marked by ‘0.5’ in the third column. The fourth column of Table 1 indicates the presence (by ‘1’) or not (by ‘0’) of an equatorial outflow or a ring/torus; a number ‘0.5’ indicates likely equatorial mass concentration.
@@
NGC 1360 142 [Mi18n] 1 [GD08] 0
MyCn 18 18.15 [Mi18m] 1 [Oc00] 0
NGC 2346 16.00 [Br19] 0.5 [Go19] 1
Sp 3 4.81 [Mi19s] 1 [Mi19s] 0.5
IC 4776 3.11 [Ni19i] 1 [So17] 0.5
NGC 7293 2.77 [Al20] 1 [Me13] 0
IC 2149 2.63877 [Al24] 1 [Va02] 1
IC 4593 2.50369 [Al21] 1 [To20] 0
NGC 2392 1.90 [Mi19n] 1 [Gu21] 0
FG 1 1.195 [We18] 1 [Lo93] 0
Necklace 1.16 [Mi13] 1[Co11] 1
Hen 2-283 1.128 [We18] 0.5 [We18] 0.5
Hen 2-161 $\simeq 1$ [Jo15] 1[Sa11] 1
Abell 65 1 [Hi15] 0.5 [Hu13] 0
HaWe 6 0.99 [Bh24] 0.5 [Zu24] 0
K 1-2 0.676 [Ex03] 1 [Co99] 0
M 2-19 0.670 [Mi08] 0.5 [Mi08] 1
M2-11 0.609 [Jo14] 1 [Jo14] 1
@@
Note: 36 PNe with binary central stars and high-quality images indicating jets. The first column is the name of the PN, and the second is the orbital period of the central star and the source for it, as appears in the site built and maintained by David (e.g., Jones 2024). The third column indicates robust morphological indications for jets (marked by ‘1’) or likely morphological indications for jets (‘0.5’), with the source of the image I used. The fourth column indicates robust (‘1’), likely (‘0.5’), or no (‘0’) morphological indications for an equatorial outflow, i.e., disk, torus, or ring.
@@
Abbreviation: P: orbital period; PN: planetary nebula; R: ring (including torus or a disk)
@@
References: Al24: Aller et al. (2024); Al20: Aller et al. (2020); Bh24; Bhattacharjee et al. (2024); Br19: Brown et al. (2019); Co99: Corradi et al. (1999): Co11: Corradi et al. (2011); Co14: Corradi et al. (2014); Co15: Corradi et al. (2015); Ex03: Exter et al. (2003); GD08: García-Díaz et al. (2008); Go19: Gómez-Muñoz et al. (2019); Gu21: Guerrero et al. (2021); GM12: Guerrero & Miranda (2012); Hi15: Hillwig et al. (2015); Hi22: Hillwig et al. (2022); Hu13: Huckvale et al. (2013); Jo10: Jones et al. (2010); Jo14: Jones et al. (2014); Jo15: Jones et al. (2015); Jo16: Jones et al. (2016); Jo19: Jones et al. (2019); Jo20: Jones et al. (2020); Lo93: Lopez et al. (1993); Ma96: Manchado et al. (1996); Me13: Meaburn et al. (2013); Mt07: Mitchell et al. (2007); Mi08: Miszalski et al. (2008); Mi09: Miszalski et al. (2009); Mi11e: Miszalski et al. (2011a); Mi11j: Miszalski et al. (2011c); Mi11a: Miszalski et al. (2011b); Mi13: Miszalski et al. (2013); Mi18n: Miszalski et al. (2018a); Mi18m: Miszalski et al. (2018b); Mi19s: Miszalski et al. (2019a); Mi19n: Miszalski et al. (2019b); Oc00: O’Connor et al. (2000); Sa11: Sahai et al. (2011); Sc92: Schwarz et al. (1992); SG15: Santander-García et al. (2015); So17: Sowicka et al. (2017); To20: Toalá et al. (2020); Va02: Vázquez et al. (2002); We18: Wesson et al. (2018); Zu24: https://brentenriegel.at/ergebnisse/a240103-btb-hawe-8-planetarischer-nebel
@@
Specifically, there are cases of directly observed jets, e.g., ETHOS 1 and the Necklace, while in some cases that have no observable jets, the bipolar structure robustly indicates jets, like the couple of misaligned lobes of M 2-26 and the extended lobes of Ou 5. In some cases, the bipolar structure might be attributed to the collimation of fast wind by an equatorial dense gas; in those cases, I marked in the third column of Table 1 the number 0.5 (although I think these are also shaped by jets).
@@ 3 Implications for CEE
However, I expect the jetted-to-ring ratio due to the CEE interaction itself to be much larger for two reasons. (1) Equatorial outflows from CEE are dense and slow and stay bright even as they expand. On the other hand, jets expand fast and disperse on a shorter time scale. I expect that a significant fraction of PNe that presently have no morphological signatures of jets were shaped by jets. (2) In some PNe, I expect jets to shape an equatorial dense outflow rather than the direct CEE mass ejection. Akashi et al. (2015) demonstrated the formation of such rings with three-dimensional hydrodynamical simulations. They present the equatorial ring of SN 1987a and the Necklace as specific examples. I marked the Necklace as having robust signatures of jets and a ring. However, the claim of Akashi et al. (2015) is that the jets shaped the ring rather than direct CEE mass ejection. I estimate that the jetted-to-ring ratio due to the CEE mass ejection is much larger than the observed one. I crudely estimate this ratio to be twice the observed one or more
$Q_{\rm JR,CEE}\gtrsim 3.$ (2)
The determination of this ratio is for the future, as well as the inclusion of other processes that influence jets and equatorial mass outflows, like magnetic fields (e.g., Nordhaus et al. 2007). At this time, I conclude that jets are more common than dense equatorial outflows concerning the CEE process.
@@ 3.2 The jets’ source
Some hydrodynamic simulations of the CEE obtain a funnel in the common envelope along the angular momentum axis of the binary system, a funnel that might launch jets (e.g., Chamandy et al. 2018, 2020; Zou et al. 2020, 2022; Moreno et al. 2022; Ondratschek et al. 2022; Vetter et al. 2024). Based on analytical calculations, the launch of jets from the common envelope funnel was suggested many years ago (Soker, 1992). However, the common envelope is not expected to precess and hence cannot form a point-symmetrical PN. Jets can change directions due to instabilities, but the two opposite jets change direction uncorrelated, hence do not form a point-symmetrical morphology (e.g., Ondratschek et al. 2022). Moreover, the jets that Ondratschek et al. (2022) find in their simulations are too weak to explain the shaping of PNe, i.e., they bent close to the envelope.
Blackman & Lucchini (2014) who examined the kinetic properties of axial-symmetric outflows from 19 pre-PNe, which were observed and analyzed by Bujarrabal et al. (2001) and Sahai et al. (2008), concluded that main sequence companions to the AGB progenitors of the pre-PNe should accrete at a very high rate to explain these kinetic properties. Accretion might in some pre-PNe be via a Roche lobe overflow (RLOF), or, more likely, for most or all pre-PNe, during a CEE and sometimes during a GEE. There are indications that the main sequence companion accretes mass during the CEE and/or shortly before the CEE, namely, during the GEE. An example is the larger radius than that of a main sequence star with an equal mass of the companion in Hen 2-155, which probably results from rapid mass accretion (Jones et al., 2015).
The relation of the shaping jets to the binary system is evident from the general alignment of the angular momentum axes of the central binaries of PNe with the PNe main symmetry axes (Hillwig et al., 2016).
@@ 3.3 The relation of the jets to CEE
Motivated by the present finding that jet-launching is the most robust observable of the CEE and cannot be ignored, I consider the possible implications to the energy budget. The launching of jets requires the main sequence companion to accrete mass. The jets allow this accretion (e.g., Shiber et al. 2016) operating in a valve mechanism that releases pressure (Chamandy et al., 2018). Applying the virial theorem, the accretion process onto the main sequence companion releases energy of
@@
where $M_{\rm MS}$ and $R_{\rm MS}$ are the companion’s initial mass and average radius during the accretion process, and $M_{\rm acc}$ is the accreted mass. The released orbital energy, if the leftover envelope mass is low, when the orbital radius is $a$, is
@@
$a\gtrsim\frac{M_{\rm core}}{M_{\rm acc}}R_{\rm MS}=5\left(\frac{M_{\rm core}}{10M_{\rm acc}}\right)\left(\frac{R_{\rm MS}}{0.5R_{\odot}}\right)R_{\odot}.$ (5)
Typical values can be $M_{\rm MS}=0.2-0.5M_{\odot}$, $M_{\rm core}=0.6M_{\odot}$, and $M_{\rm acc}\approx 0.06M_{\odot}$. Jets might carry a fraction of 0.1-0.2 of the accretion energy. The rest is radiated away. Not all the energy deposited by the different processes (accretion, orbital, recombination) ends in unbinding the envelope; the convection can efficiently carry energy out where it is radiated away (e.g., Sabach et al. 2017; Wilson & Nordhaus 2019, 2020, 2022; Noughani et al. 2024). I do not discuss recombination energy here (for a recent paper and references, see Chamandy et al. 2024; recombination seems to play a small role in envelope ejection).

### v2

@@ 1 Introduction
There are more than a hundred bipolar and elliptical PNe, including some post-asymptotic giant branch nebulae and pre-PNe, having a central binary system (e.g., Miszalski et al. 2019c; Orosz et al. 2017; Jones 2020, 2024). The short orbital periods of these binary systems show that they have experienced a common envelope evolution (CEE) phase. In post-AGB systems and PNe with long orbital periods, months to years (e.g., Van Winckel et al. 2014), the systems might have experienced the grazing envelope evolution (GEE). The morphologies of many of these suggest that jets shaped these PNe, pointing at the direct link between CEE and jet-launching. In this study, I quantify this occurrence rate and compare it to the occurrence rate of equatorial outflows (ring/torus/disk).
@@ 2 The sample of post-CEE PNe with jets
I examined the catalog of PNe with a binary central star built and maintained by David Jones [^1] https://www.drdjones.net/bCSPN/ (Jones & Boffin 2017; Boffin & Jones 2019; Jones 2024). I do not consider the uncertain cases in that catalog (marked by * or ** there). This leaves 108 PNe. I could find high-quality images to decide whether jets exist in 50 PNe. Out of these 50 PNe, I saw clear indications for jets or morphologies robustly suggesting jets in 28 PNe, marked by the number ‘1’ in the third column of Table 1. In 8 PNe, I found morphologies that likely suggest jets, marked by ‘0.5’ in the third column. The fourth column of Table 1 indicates the presence (by ‘1’) or not (by ‘0’) of an equatorial outflow or a ring/torus; a number ‘0.5’ indicates likely equatorial mass concentration. A subscript ‘j’ near the reference indicates that this reference suggests jets (with possible references to earlier claims for jets). An up-script points to another reference that claims for jets. Otherwise, the claim for jets is in this study. The subscript ‘r’ similarly refers to the evidence for a ring (equatorial outflow).
@@
NGC 1360 142 [Mi18n] 1 ${}^{\rm[GD08]_{\rm j}}$ 0
MyCn 18 18.15 [Mi18m] 1 ${}^{\rm[Oc00]_{\rm j}}$ 0
NGC 2346 16.00 ${}^{\rm[Br19]_{\rm r}}$ 0.5 [Go19] 1
Sp 3 4.81 [Mi19s] 1 ${}^{\rm[Mi19s]_{\rm j,r}}$ 0.5
IC 4776 3.11 [Mi19i] 1 ${}^{\rm[So17]_{\rm j}}$ 0.5
NGC 7293 2.77 [Al20] 1 ${}^{\rm[Me13]_{\rm j}}$ 0
IC 2149 2.63877 ${}^{\rm[Al24]_{\rm r}}$ 1 ${}^{\rm[Va02]^{\#1}}$ 1
IC 4593 2.50369 [Al21] 1 ${}^{\rm[To20]_{\rm j}}$ 0
NGC 2392 1.90 [Mi19n] 1 ${}^{\rm[Gu21]_{\rm j}}$ 0
Fg 1 1.195 [We18] 1 ${}^{\rm[Lo93]_{\rm j}}$ 0
Necklace 1.16 ${}^{\rm[Mi13]_{\rm j}}$ 1${}^{\rm[Co11]_{\rm r}}$ 1
Hen 2-283 1.15 [We18] 0.5 [We18] 0.5
Hen 2-161 $\simeq 1$ ${}^{\rm[Jo15]_{\rm j}}$ 1${}^{\rm[Sa11]_{\rm r}}$ 1
Abell 65 1 [Hi15] 0.5 ${}^{\rm[Hu13]_{\rm j}}$ 0
HaWe 8 0.99 [Bh24] 0.5 [Zu24] 0
K 1-2 0.676 [Ex03] 1 ${}^{\rm[Co99]_{\rm j}}$ 0
M 2-19 0.670 [Mi08] 0.5 ${}^{\rm[Mi08]_{\rm r}}$ 1
Hen 2-11 0.609 [Jo14] 1 ${}^{\rm[Jo14]_{\rm j}}$ 1
@@
Note: 36 PNe with binary central stars and high-quality images indicating jets. The first column is the name of the PN, and the second is the orbital period of the central star and the source for it, as appears in the site built and maintained by David Jones (e.g., Jones 2024). The third column indicates robust morphological indications for jets (marked by ‘1’) or likely morphological indications for jets (‘0.5’), with the source of the image I used. A subscript ‘j’ in one of the two references in a row indicates that the binary or the image reference claims for jets, while an up-script mentions another reference to support jets. Otherwise, the claim for jets is in this study. The same holds for a subscript ‘r’ for the indication of an equatorial dense gas (ring). The fourth column indicates robust (‘1’), likely (‘0.5’), or no (‘0’) morphological indications for an equatorial outflow, i.e., disk, torus, or ring.
@@
Abbreviation: P: orbital period; PN: planetary nebula; R: ring (including torus or a disk); PN G 283: PN G283.7-05.1.
@@
References: Al24: Aller et al. (2024); Al20: Aller et al. (2020); Bh24; Bhattacharjee et al. (2024); Br19: Brown et al. (2019); Co99: Corradi et al. (1999): Co11: Corradi et al. (2011); Co14: Corradi et al. (2014); Co15: Corradi et al. (2015); Ex03: Exter et al. (2003); GD08: García-Díaz et al. (2008); Go19: Gómez-Muñoz et al. (2019); Gu21: Guerrero et al. (2021); GM12: Guerrero & Miranda (2012); Hi15: Hillwig et al. (2015); Hi22: Hillwig et al. (2022); Hu13: Huckvale et al. (2013); Jo10: Jones et al. (2010); Jo14: Jones et al. (2014); Jo15: Jones et al. (2015); Jo16: Jones et al. (2016); Jo19: Jones et al. (2019); Jo20: Jones et al. (2020); Lo93: Lopez et al. (1993); Ma96: Manchado et al. (1996); Me13: Meaburn et al. (2013); Mt07: Mitchell et al. (2007); Mi08: Miszalski et al. (2008); Mi09: Miszalski et al. (2009); Mi11e: Miszalski et al. (2011a); Mi11j: Miszalski et al. (2011c); Mi11a: Miszalski et al. (2011b); Mi13: Miszalski et al. (2013); Mi18n: Miszalski et al. (2018a); Mi18m: Miszalski et al. (2018b); Mi19s: Miszalski et al. (2019a); Mi19i: Miszalski et al. (2019c); Mi19n: Miszalski et al. (2019b); Oc00: O’Connor et al. (2000); Sa11: Sahai et al. (2011); Sc92: Schwarz et al. (1992); SG15: Santander-García et al. (2015); So17: Sowicka et al. (2017); To20: Toalá et al. (2020); Va02: Vázquez et al. (2002); We18: Wesson et al. (2018); Zu24: https://brentenriegel.at/ergebnisse/a240103-btb-hawe-8-planetarischer-nebel; #1: claim for jets in Balick et al. (1993). #2: claim for a ring in Manchado, Stanghellini, & Guerrero (1996). #3: claim for a ring in Miszalski et al. (2011c).
@@
Specifically, there are cases of directly observed jets, e.g., ETHOS 1 and the Necklace, while in some cases that have no observable jets, the bipolar structure robustly indicates jets, like the couple of misaligned lobes of M 2-46 and the extended lobes of Ou 5. In some cases, the bipolar structure might be attributed to the collimation of fast wind by an equatorial dense gas; in those cases, I marked in the third column of Table 1 the number 0.5 (although I think these are also shaped by jets). Figure 1 presents four PNe that demonstrate some of the classifications.
Figure 1: Four PNe demonstrating jets and equatorial outflows. (a) An image of M 2-19 adapted from Miszalski et al. (2008). (b) An HST image of NGC 2346. In panels (a) and (b) I marked the equatorial outflow and added the dashed line in each panel to indicate the polar axis. These two PNe demonstrate equatorial mass concentration, hence the value ‘1’ in the fourth column of Table 1.There are no direct indications of jets, only indirect indications by the openings to the two polar directions, hence the number ‘0.5’ in the third column. (c) An image of The Necklace adapted from Corradi et al. (2011) clearly demonstrates the existence of jets and an equatorial ring. (d) An image of NGC 2392 from Guerrero et al. (2021) demonstrating that, in some cases, a deep analysis is required to reveal the jets. The blue regions are the approaching jets at $-190{~{}\rm km}{~{}\rm s}^{-1}$ to $-150{~{}\rm km}{~{}\rm s}^{-1}$, while the red regions are the receding jet $+150{~{}\rm km}{~{}\rm s}^{-1}$ to $+190{~{}\rm km}{~{}\rm s}^{-1}$. The white regions near the center are the nebula at the velocity interval of $-135{~{}\rm km}{~{}\rm s}^{-1}$ to $+135{~{}\rm km}{~{}\rm s}^{-1}$. There is no indication of a dense equatorial outflow.
@@ 3.1 Jetted-to-ring post-CEE PNe ratio
However, I expect the jetted-to-ring ratio due to the CEE interaction itself to be much larger for two reasons. (1) Detection biases. (1.1) Equatorial outflows from CEE are dense and slow and stay bright even as they expand. On the other hand, jets expand fast and disperse on a shorter time scale. One example is the Necklace (panel c of Figure 1). The jets are fainter and at a much larger distance from the center than the ring. A similar PN but much older, by thousands of years, might reveal only its ring and not the jets. (1.2) A ring is bright and prominent from all directions. Jets are hard to detect when pointing at the observer. The PN NGC 2392 (panel c of Figure 1) is an example. Only a deep analysis with Doppler shift reveals its jets. (2) In some PNe, I expect jets to shape an equatorial dense outflow rather than the direct CEE mass ejection. Akashi et al. (2015) and Shiber (2018) demonstrated the formation of such rings with three-dimensional hydrodynamical simulations. Akashi et al. (2015) present the equatorial ring of SN 1987a and the Necklace as specific examples. I marked the Necklace as having robust signatures of jets and a ring. However, the claim of Akashi et al. (2015) is that the jets shaped the ring rather than direct CEE mass ejection. For the above reasons, I expect that jets have shaped a significant fraction of PNe that currently lack jet morphological signatures. A quantitative study is for the future, with deeper analysis of more PNe. From the large number of PNe with barely observed jets (like the Necklace and NFC 2392), I crudely estimate that the jetted-to-ring ratio due to the CEE mass ejection is approximately twice as given from the inspection of images.
$Q_{\rm JR,CEE}\approx 2\times Q_{\rm JR,obs}\gtrsim 3.$ (2)
The determination of this ratio is for the future, as well as the inclusion of other processes that influence jets and equatorial mass outflows, like magnetic fields (e.g., Nordhaus et al. 2007). I note that the possibility of missing an equatorial ring because it is not resolved is unlikely. Such a ring (torus) is very bright because it is dense and close to the center. The bright waist of Hen 2-84 (e.g., Hillwig et al. 2022) is an example. At this time, I conclude that jets are more common than dense equatorial outflows concerning the CEE process.
@@ 3.2 The jets’ source
Some hydrodynamic simulations of the CEE obtain a funnel in the common envelope along the angular momentum axis of the binary system, a funnel that might launch jets (e.g., Chamandy et al. 2018, 2020; Zou et al. 2020, 2022; Moreno et al. 2022; Ondratschek et al. 2022; Vetter et al. 2024; Gagnier & Pejcha 2024). Based on analytical calculations, the launch of jets from the common envelope funnel was suggested many years ago (Soker 1992). However, the common envelope is not expected to precess and hence cannot form a point-symmetrical PN. Jets can change directions due to instabilities, but the two opposite jets change direction uncorrelated, hence do not form a point-symmetrical morphology (e.g., Ondratschek et al. 2022). Moreover, the jets that, e.g., Ondratschek et al. (2022) and Gagnier & Pejcha (2024), find in their simulations are too weak to explain the shaping of PNe, i.e., they bent close to the envelope.
Blackman & Lucchini (2014) who examined the kinetic properties of axial-symmetric outflows from 19 pre-PNe, which were observed and analyzed by Bujarrabal et al. (2001) and Sahai et al. (2008), concluded that main sequence companions to the AGB progenitors of the pre-PNe should accrete at a very high rate to explain these kinetic properties. Accretion might in some pre-PNe be via a Roche lobe overflow (RLOF), or, more likely, for most or all pre-PNe, during a CEE and sometimes during a GEE. There are indications that the main sequence companion accretes mass during the CEE and/or shortly before the CEE, namely, during the GEE. An example is the larger radius than that of a main sequence star with an equal mass of the companion in Hen 2-155, which probably results from rapid mass accretion (Jones et al. 2015).
The relation of the shaping jets to the binary system is evident from the general alignment of the angular momentum axes of the central binaries of PNe with the PNe main symmetry axes (Hillwig et al. 2016).
@@ 3.3 The relation of the jets to CEE
Motivated by the present finding that jet-launching is the most robust observable of the CEE and cannot be ignored, I consider the possible implications to the energy budget. The launching of jets requires the main sequence companion to accrete mass. The jets allow this accretion (e.g., Shiber et al. 2016) operating in a valve mechanism that releases pressure (Chamandy et al. 2018). Applying the virial theorem, the accretion process onto the main sequence companion releases energy of
@@
where $M_{\rm MS}$ and $R_{\rm MS}$ are the companion’s initial mass and average radius during the accretion process, and $M_{\rm acc}$ is the accreted mass. Recent studies suggest that main sequence stars can accrete mass at high rates with only minor expansion if jets remove the outskirts of the envelope (Bear & Soker 2025; Scolnic et al. 2025). The released orbital energy, if the leftover envelope mass is low, when the orbital radius is $a$, is
@@
$a\gtrsim\frac{M_{\rm core}}{M_{\rm acc}}R_{\rm MS}=10\left(\frac{M_{\rm core}}{20M_{\rm acc}}\right)\left(\frac{R_{\rm MS}}{0.5R_{\odot}}\right)R_{\odot}.$ (5)
Typical values can be $M_{\rm MS}=0.15-0.4M_{\odot}$, $R_{\rm MS}=0.4-0.7R_{\odot}$, and $M_{\rm core}=0.5-0.7M_{\odot}$ (e.g., Hillwig et al. 2010; De Marco et al. 2011; Jones et al. 2015). I scale with a value of $M_{\rm acc}\approx 0.03M_{\odot}$. Jets might carry a fraction of 0.1-0.2 of the accretion energy. The rest is radiated away. Not all the energy deposited by the different processes (accretion, orbital, recombination) ends in unbinding the envelope; the convection can efficiently carry energy out where it is radiated away (e.g., Sabach et al. 2017; Wilson & Nordhaus 2019, 2020, 2022; Noughani et al. 2024). I do not discuss recombination energy here (for a recent paper and references, see Chamandy et al. 2024; recombination seems to play a small role in envelope ejection).